If k is a positive integer, what is the average of 3^k and 3^(k+2)?
Answer by Alan3354(69443)

If k is a positive integer, what is the average of 3^k and 3^(k+2)?
-----------
Avg = (3^k + 3^(k+2))/2 = (3^k + 9*3^k)/2
Avg = 5*3^k